# Annotation-Efficient Universal Honesty Alignment

### v1

@@ 2.1 Training-free Confidence Investigation
Early works linked confidence to token probabilities (Guo et al., 2017; Desai & Durrett, 2020; Jiang et al., 2021), but these signals are often miscalibrated in free-form generation where probabilities can be dominated by semantically irrelevant tokens. To address this, self-consistency-based methods measure confidence from the semantic consistency of multiple generations (Manakul et al., 2023), achieving the most reliable results among training-free methods. Another line explores verbalized confidence, where LLMs explicitly their confidence in words (Lin et al., 2022; Yin et al., 2023; Tian et al., 2023), though these models often remain overconfident.
@@ 2.2 Training-based Confidence Calibration
These studies leverage correctness annotations to calibrate model confidence, achieving better performance than training-free methods, and can be broadly divided into two categories. One line leverages LLMs’ internal states to predict confidence either after or even before generation (Azaria & Mitchell, 2023; Chen et al., 2024; Wang et al., 2024). Another line trains models to verbalize confidence reliably (Lin et al., 2022; Zhang et al., 2024). All these methods rely on correctness annotations, and achieving optimal performance requires high annotation costs. Although some works (Zhang et al., 2024; Tjandra et al., 2024) exploit LLMs’ internal uncertainty as a supervision signal, it is only used to determine abstention rather than to teach models to express their own confidence. Apart from that, all the above methods are trained only on small-scale datasets.
@@ 5 HonestyBench
HonestyBench-Eval provides evaluation across both in-domain and OOD scenarios. In-domain evaluation uses the test or development splits of the five datasets included in HonestyBench-Train. Out-of-domain (OOD) evaluation covers five additional factual QA datasets—SQuAD (Rajpurkar et al., 2016), WebQuestions (WQ) (Berant et al., 2013), ComplexWebQuestions (CWQ) (Talmor & Berant, 2018), MuSiQue (Trivedi et al., 2022), and PopQA (Mallen et al., 2022)—spanning single-hop, multi-hop, and template-generated questions in diverse domains. The in-domain evaluation contains approximately 38k QA pairs, while the out-of-domain evaluation contains approximately 33k QA pairs. As in training, each question is annotated with both consistency and correctness scores.
@@ 6 Experimental Setup
For QA performance, we measure accuracy by verifying whether the model’s greedy search output matches any ground-truth answer using Qwen2.5-32B-Instruct (scored as 1 if correct, 0 if incorrect). To evaluate honesty alignment, we adopt the widely used AUROC (Hanley & McNeil, 1982) (Area Under the Receiver Operating Characteristic Curve) metric. AUROC measures a model’s ability to distinguish correct from incorrect predictions: higher values indicate that the model assigns higher confidence to correct answers. It is computed as the area under the curve plotting the true positive rate against the false positive rate at varying confidence thresholds. A value of 1 represents perfect discrimination, while 0.5 corresponds to random guessing. We also evaluate honesty alignment using ECE (Guo et al., 2017) in §B.
@@
We compare EliCal with six representative training-free baselines and two training-based baselines. The training-free methods include three types, each with two variants: 1) Probabilistic confidence (Prob): sequence-level generation probability, with length-normalized version (N-Prob); 2) Self-consistency (Consis) (Manakul et al., 2023; Ho et al., 2020): measured via lexical similarity (Consis-Lex) or an LLM for semantic similarity (Consis-Sem); 3) Verbalized confidence (Verbal) (Xiong et al., 2023): model expresses confidence in natural language, in zero-shot (Verbal-0) and few-shot (Verbal-10) settings. The training-based baselines are: 1) Elicitation-Only (Eli-Only): learning from Consis-Sem, and 2) Calibration-Only (Cal-Only) (Yang et al., 2023): learning from correctness from scratch. Implementation details are in §E.
@@ 7 Results and Analysis
Training-based Eli-Only 77.86 86.23 77.27 54.36 62.05 71.19 60.66 76.61 74.77 66.56 74.60 69.66
@@ 7.1 Ablation
8 Conclusion
In this paper, we propose EliCal, an annotation-efficient two-stage training framework for honesty alignment, and introduce HonestyBench, a large-scale benchmark enabling universal honesty training and comprehensive evaluation. Our results demonstrate that EliCal significantly improves model confidence expression with minimal labeled data, while HonestyBench supports the development of models that excel across diverse tasks. This work sets the stage for scalable, high-performance, and data-efficient honesty alignment in real-world AI applications.
@@ A.1 Training-free Confidence Investigation
1) Probability-based Confidence. A common approach links model confidence to the probabilities assigned during token generation (Guo et al., 2017; Desai & Durrett, 2020; Jiang et al., 2021; Kadavath et al., 2022; Si et al., 2022; Kuhn et al., 2023). Early work (Guo et al., 2017) revealed that modern neural networks such as ResNet (He et al., 2016) tend to produce overconfident predictions, and introduced temperature scaling as a correction. Later, Desai & Durrett (2020) showed that pretrained language models such as BERT (Devlin, 2018) achieve more reliable calibration compared to models without pretraining. As generative models became prominent, Jiang et al. (2021) reported that T5 (Raffel et al., 2020) still exhibited miscalibration, often being more confident than warranted. Recent studies highlight that LLMs appear well calibrated in structured tasks (e.g., multiple-choice QA) under suitable prompting (Kadavath et al., 2022; Si et al., 2022), but their probabilities deviate substantially from correctness in free-form generation.
@@ A.2 Training-based Confidence Calibration
A more recent stream of research investigates whether the internal representations of LLMs encode signals about factual correctness (Azaria & Mitchell, 2023; Su et al., 2024; Chen et al., 2024; Wang et al., 2024; Ni et al., 2025). Azaria & Mitchell (2023) showed that hidden states can reflect factuality judgments. Building on this, Su et al. (2024) and Chen et al. (2024) found that post-generation activations capture whether a model’s own outputs are factual. More recently, Wang et al. (2024); Ni et al. (2025) demonstrated that pre-generation states already carry predictive cues, enabling estimation of correctness before the answer is fully produced.
@@ Appendix B Further Analysis Using ECE
In addition to evaluating whether the model’s confidence can distinguish between questions it can and cannot answer correctly, we also hope that the confidence values themselves are meaningful, i.e., that confidence reflects accuracy. This is also why we do not consider entropy-based methods, since the value of entropy itself does not represent a confidence score between 0 and 1, and its value does not provide a direct characterization of ability. We use ECE (Expected Calibration Error) to measure this, which can be formulated as:
@@
where the predictions are partitioned into $M=10$ bins according to their confidence scores, $B_{m}$ denotes the set of samples in the $m$-th bin, and $n$ is the total number of samples. For each bin, $\text{acc}(B_{m})$ is the empirical accuracy of the predictions and $\text{conf}(B_{m})$ is their average confidence. The absolute difference $\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|$ quantifies the miscalibration in that bin, and the overall ECE is obtained as the sample-size–weighted average across bins. As shown in Figure 11, EliCal and Cal-Only achieve similarly low ECE in most cases, indicating that both learn calibrated confidence overall. However, when labeled data is limited (Figure 5), Cal-Only shows worse AUROC, suggesting that it captures only global trends (e.g., confidence close to average accuracy) but lacks fine-grained discriminative ability.
@@ Appendix C Implementation Details
We use Qwen2.5-32B-Instruct to measure consistency between two responses (See Figure 12). Yang et al. (2023) show that full-parameter fine-tuning for honesty alignment can negatively impact the model’s QA performance. To avoid affecting the model’s original capabilities, we train with LoRA (Hu et al., 2022) and a linear head to output a confidence score, where we set rank=8 and $\alpha$=16. We use AdamW (Loshchilov & Hutter, 2017) as the optimizer, MSE (Mean Square Error) as the loss function and conduct training with the SFTTrainer from trl [^1] https://huggingface.co/docs/trl/sft_trainer, using a batch size of 16 and accumulation steps of 8. For answer generation, we use vLLM [^2] https://docs.vllm.ai/en/latest/. For each question, we generate one greedy answer with temperature = 0, and sample 20 answers with temperature = 1, top-p = 0.95, and top-k = 50. Checkpoints for all training methods are selected using the in-domain test set. All other parameters are kept at their default settings. All the prompts can be seen in §G.
@@ Appendix E Details of Baselines
In this section, we describe how each training-free baseline method is implemented. For the question $q$, suppose the greedy answer generated by the model is $\tilde{r}$ and the set of sampled answers is $\hat{\mathcal{R}}$. $\hat{\mathcal{R}}$ contains 20 responses in our paper. Using the token generation probabilities of the model to represent confidence is a common approach (Guo et al., 2017; Desai & Durrett, 2020; Jiang et al., 2021; Ni et al., 2024b); in this work, we implement two versions.
Prob

### v2

@@ 2.1 Training-free Confidence Investigation
Early works linked confidence to token probabilities (Guo et al., 2017; Desai and Durrett, 2020; Jiang et al., 2021), but these signals are often miscalibrated in free-form generation where probabilities can be dominated by semantically irrelevant tokens. To address this, self-consistency-based methods measure confidence from the semantic consistency of multiple generations (Manakul et al., 2023), achieving the most reliable results among training-free methods. Another line explores verbalized confidence, where LLMs explicitly their confidence in words (Lin et al., 2022; Yin et al., 2023; Tian et al., 2023), though these models often remain overconfident.
@@ 2.2 Training-based Confidence Calibration
These studies leverage correctness annotations to calibrate model confidence, achieving better performance than training-free methods, and can be broadly divided into two categories. One line leverages LLMs’ internal states to predict confidence either after or even before generation (Azaria and Mitchell, 2023; Chen et al., 2024; Wang et al., 2024). Another line trains models to verbalize confidence reliably (Lin et al., 2022; Zhang et al., 2024). All these methods rely on correctness annotations, and achieving optimal performance requires high annotation costs. Although some works (Zhang et al., 2024; Tjandra et al., 2024) exploit LLMs’ internal uncertainty as a supervision signal, it is only used to determine abstention rather than to teach models to express their own confidence. Apart from that, all the above methods are trained only on small-scale datasets.
@@ 5 HonestyBench
HonestyBench-Eval provides evaluation across both in-domain and OOD scenarios. In-domain evaluation uses the test or development splits of the five datasets included in HonestyBench-Train. Out-of-domain (OOD) evaluation covers five additional factual QA datasets—SQuAD (Rajpurkar et al., 2016), WebQuestions (WQ) (Berant et al., 2013), ComplexWebQuestions (CWQ) (Talmor and Berant, 2018), MuSiQue (Trivedi et al., 2022), and PopQA (Mallen et al., 2022)—spanning single-hop, multi-hop, and template-generated questions in diverse domains. The in-domain evaluation contains approximately 38k QA pairs, while the out-of-domain evaluation contains approximately 33k QA pairs. As in training, each question is annotated with both consistency and correctness scores.
@@ 6 Experimental Setup
For QA performance, we measure accuracy by verifying whether the model’s greedy search output matches any ground-truth answer using Qwen2.5-32B-Instruct (scored as 1 if correct, 0 if incorrect). To evaluate honesty alignment, we adopt the widely used AUROC (Hanley and McNeil, 1982) (Area Under the Receiver Operating Characteristic Curve) metric. AUROC measures a model’s ability to distinguish correct from incorrect predictions: higher values indicate that the model assigns higher confidence to correct answers. It is computed as the area under the curve plotting the true positive rate against the false positive rate at varying confidence thresholds. A value of 1 represents perfect discrimination, while 0.5 corresponds to random guessing. We also evaluate honesty alignment using ECE (Guo et al., 2017) in §B.
@@
We compare EliCal with six representative training-free baselines and two training-based baselines. The training-free methods include three types, each with two variants: 1) Probabilistic confidence (Prob): sequence-level generation probability, with length-normalized version (N-Prob); 2) Self-consistency (Consis) (Manakul et al., 2023; Ho et al., 2020): measured via lexical similarity (Consis-Lex) or an LLM for semantic similarity (Consis-Sem); 3) Verbalized confidence (Verbal) (Xiong et al., 2023): model expresses confidence in natural language, in zero-shot (Verbal-0) and few-shot (Verbal-10) settings. The training-based baselines are: 1) Elicitation-Only (Eli-Only): learning from Consis-Sem, and 2) Calibration-Only (Cal-Only) (Yang et al., 2023): learning from correctness from scratch. We also include two recent temperature-scaling-based methods Thermometer (Shen et al., 2024) and DACA (Luo et al., 2025). Implementation details are in §E.
@@ 7 Results and Analysis
Training-based Thermometer 58.15 63.38 58.08 55.24 67.58 59.48 60.23 62.90 56.98 61.90 68.76 63.88
DACA 61.69 72.54 66.39 65.03 71.06 67.70 66.79 65.71 62.62 73.98 76.40 70.98
@@ 7.1 Ablation
Effects of sample size $k$.
An ablation study is conducted to investigate the effect of the number of samples $k$ used to compute self-consistency–based confidence with Qwen2.5-7B-Instruct. In Stage 1, different values of $k\in\{2,5,10,20\}$ are evaluated and the results are shown in Table 3. We observe that for Consis-Sem, the estimation quality consistently improves as $k$ increases. In contrast, Eli-Only exhibits only minor performance variation across different values of $k$. Notably, even at $k=2$, although Consis-Sem introduces additional noise, the resulting signal remains sufficiently informative, enabling the model to learn to express its internal confidence from large-scale unlabeled data. Moreover, EliCal demonstrates highly stable performance across different values of $k$, indicating strong robustness to the choice of sampling count.
Table 3: AUROC scores of different methods across varying sample sizes $k$ on Qwen2.5-7B-Instruct. Bold indicates the best result on each dataset.
K Methods In-Domain Evaluation OOD Evaluation
NQ TQ HQ 2Wiki Pararel Avg. Squad WQ CWQ MSQ PopQA Avg.
2 Consis-Sem 75.42 83.04 73.13 51.13 60.93 68.04 62.39 70.52 70.94 63.55 66.26 65.58
Eli-Only 78.46 86.03 77.30 52.48 63.59 70.70 60.61 77.20 75.24 64.67 75.32 69.90
EliCal (1k) 82.37 87.32 84.60 82.13 85.04 84.41 79.03 81.04 79.72 78.45 91.94 84.80
5 Consis-Sem 78.72 87.76 77.85 53.14 62.05 71.44 64.59 73.88 74.99 67.63 68.69 68.29
Eli-Only 78.86 86.09 77.82 54.10 65.45 71.54 62.34 77.42 76.12 67.13 77.92 71.88
EliCal (1k) 82.49 87.46 84.46 82.37 84.98 84.52 79.05 80.60 79.17 78.70 91.95 84.74
10 Consis-Sem 79.70 89.40 79.41 54.41 62.47 72.78 65.68 75.25 76.58 69.70 69.82 69.54
Eli-Only 78.83 86.26 77.36 54.17 63.41 71.36 61.51 77.61 75.58 66.52 76.05 70.71
EliCal (1k) 82.58 87.56 84.64 82.22 84.59 84.51 78.88 80.87 79.28 78.05 91.98 84.68
20 Consis-Sem 80.68 90.20 80.12 55.40 62.93 73.62 66.16 76.26 77.50 70.74 70.44 70.20
EliCal (1k) 82.38 87.51 84.48 82.05 84.31 84.36 78.48 80.11 79.85 78.09 91.74 84.47
Scalability of EliCal.
To investigate the scalability of EliCal, we conduct experiments with Qwen2.5-32B-Instruct on HonestyBench, and the results are reported in Table 4. We observe that EliCal (1k) still significantly outperforms Cal-Only (1k), indicating that our method remains effective when applied to larger models. These findings suggest that EliCal scales well with model size.
Table 4: AUROC scores on Qwen2.5-32B-Instruct. Bold indicates the best result on each dataset.
Methods In-Domain Evaluation OOD Evaluation
NQ TQ HQ 2Wiki Pararel Avg. Squad WQ CWQ MSQ PopQA Avg.
Consis-Sem 80.18 89.07 80.78 56.62 73.11 74.57 66.89 74.97 76.24 71.31 74.66 72.10
Eli-Only 78.26 87.31 79.14 52.55 73.98 72.26 64.64 75.86 75.34 63.67 79.55 72.90
Cal-Only (1k) 75.15 81.12 79.63 78.37 80.67 79.31 73.64 72.99 76.69 69.27 87.47 79.62
[…]
Cal-Only (560k) 85.19 89.59 86.18 85.17 88.52 86.95 82.38 79.22 81.78 81.91 91.32 85.97
EliCal (560k) 85.11 89.91 86.26 85.22 89.05 87.12 82.86 80.38 82.41 83.04 91.55 86.45
8 Conclusion and Future Work
In this paper, we propose EliCal, an annotation-efficient two-stage training framework for honesty alignment, and introduce HonestyBench, a large-scale benchmark enabling universal honesty training and comprehensive evaluation. Our results demonstrate that EliCal significantly improves model confidence expression with minimal labeled data, while HonestyBench supports the development of models that excel across diverse tasks. This work sets the stage for scalable, high-performance, and data-efficient honesty alignment in real-world AI applications. In this paper, we primarily focus on single-turn, text-only, free-form QA tasks. Extending EliCal to multi-turn interactions, multimodal settings, and a broader range of task types remains an important direction for future research.
@@ 8 Conclusion and Future Work
ACKNOWLEDGEMENTS
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No. 62302486 and No. 62441229, the Innovation Project of ICT CAS under Grants No. E361140, the CAS Special Research Assistant Funding Project, and the project under Grants No. JCKY2022130C039.
@@ A.1 Training-free Confidence Investigation
1) Probability-based Confidence. A common approach links model confidence to the probabilities assigned during token generation (Guo et al., 2017; Desai and Durrett, 2020; Jiang et al., 2021; Kadavath et al., 2022; Si et al., 2022; Kuhn et al., 2023). Early work (Guo et al., 2017) revealed that modern neural networks such as ResNet (He et al., 2016) tend to produce overconfident predictions, and introduced temperature scaling as a correction. Later, Desai and Durrett (2020) showed that pretrained language models such as BERT (Devlin, 2018) achieve more reliable calibration compared to models without pretraining. As generative models became prominent, Jiang et al. (2021) reported that T5 (Raffel et al., 2020) still exhibited miscalibration, often being more confident than warranted. Recent studies highlight that LLMs appear well calibrated in structured tasks (e.g., multiple-choice QA) under suitable prompting (Kadavath et al., 2022; Si et al., 2022), but their probabilities deviate substantially from correctness in free-form generation.
@@ A.2 Training-based Confidence Calibration
A more recent stream of research investigates whether the internal representations of LLMs encode signals about factual correctness (Azaria and Mitchell, 2023; Su et al., 2024; Chen et al., 2024; Wang et al., 2024; Ni et al., 2025). Azaria and Mitchell (2023) showed that hidden states can reflect factuality judgments. Building on this, Su et al. (2024) and Chen et al. (2024) found that post-generation activations capture whether a model’s own outputs are factual. More recently, Wang et al. (2024); Ni et al. (2025) demonstrated that pre-generation states already carry predictive cues, enabling estimation of correctness before the answer is fully produced.
@@ Appendix B Further Analysis Using ECE
In addition to evaluating whether the model’s confidence can distinguish between questions it can and cannot answer correctly, we also hope that the confidence values themselves are meaningful, i.e., that confidence reflects accuracy. We use ECE (Expected Calibration Error) to measure this, which can be formulated as:
@@
where the predictions are partitioned into $M=10$ bins according to their confidence scores, $B_{m}$ denotes the set of samples in the $m$-th bin, and $n$ is the total number of samples. For each bin, $\text{acc}(B_{m})$ is the empirical accuracy of the predictions and $\text{conf}(B_{m})$ is their average confidence. The absolute difference $\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|$ quantifies the miscalibration in that bin, and the overall ECE is obtained as the sample-size–weighted average across bins. As shown in Figure 11, EliCal and Cal-Only achieve similarly low ECE in most cases, indicating that both learn calibrated confidence overall. However, when labeled data is limited (Figure 5), Cal-Only shows worse AUROC, suggesting that it captures only global trends (e.g., confidence close to average accuracy) but lacks fine-grained discriminative ability. We further compare the calibration performance of EliCal with all other baselines, and the results are presented in Table 6. It shows that among training-free methods, self-consistency–based approaches achieve relatively strong performance. Among training-based methods, in addition to EliCal, temperature-scaling–based methods also show competitive calibration results; however, their performance in terms of AUROC remains unsatisfactory (see Table 2). We find that these methods tend to assign relatively high temperatures (mostly between 5 and 10), which uniformly reduce confidence across all questions and consequently lower ECE, as the overall confidence becomes closer to the empirical accuracy. However, this strategy does not effectively improve the discriminative ability of the confidence scores to distinguish between correct and incorrect predictions. Together with the AUROC results, the ECE results further suggest that EliCal consistently achieves superior performance in both calibration and discrimination.
Table 6: ECE performance of different methods on Qwen2.5-7B-Instruct. Bold indicates the best result on each dataset.
Methods In-Domain Evaluation OOD Evaluation
NQ TQ HQ 2Wiki Pararel Avg. Squad WQ CWQ MSQ PopQA Avg.
Prob 0.39 0.53 0.32 0.31 0.46 0.40 0.31 0.54 0.33 0.13 0.16 0.25
N-Prob 0.39 0.24 0.47 0.50 0.32 0.39 0.46 0.24 0.46 0.65 0.60 0.52
Verbal-0 0.41 0.28 0.44 0.41 0.28 0.37 0.42 0.29 0.44 0.46 0.57 0.48
Verbal-10 0.27 0.23 0.26 0.15 0.30 0.22 0.35 0.32 0.36 0.46 0.36 0.36
Consis-Lex 0.07 0.07 0.16 0.17 0.03 0.12 0.14 0.07 0.13 0.32 0.26 0.20
Consis-Sem 0.13 0.04 0.11 0.28 0.25 0.16 0.31 0.15 0.16 0.21 0.29 0.27
Thermometer 0.06 0.05 0.04 0.05 0.04 0.06 0.05 0.05 0.01 0.06 0.10 0.07
DACA 0.09 0.05 0.10 0.14 0.10 0.10 0.09 0.08 0.06 0.06 0.15 0.11
Elicitation 0.09 0.03 0.06 0.24 0.21 0.13 0.31 0.09 0.11 0.19 0.24 0.24
Cal-Only(1k) 0.12 0.16 0.18 0.13 0.13 0.15 0.17 0.15 0.16 0.12 0.07 0.12
EliCal(1k) 0.05 0.06 0.03 0.05 0.08 0.05 0.10 0.06 0.05 0.08 0.02 0.06
@@ Appendix C Implementation Details
We use Qwen2.5-32B-Instruct to measure consistency between two responses (See Figure 12). Yang et al. (2023) show that full-parameter fine-tuning for honesty alignment can negatively impact the model’s QA performance. To avoid affecting the model’s original capabilities, we train with LoRA (Hu et al., 2022) and a linear head to output a confidence score, where we set rank=8 and $\alpha$=16. We use AdamW (Loshchilov and Hutter, 2017) as the optimizer, MSE (Mean Square Error) as the loss function and conduct training with the SFTTrainer from trl [^1] https://huggingface.co/docs/trl/sft_trainer, using a batch size of 16 and accumulation steps of 8. For answer generation, we use vLLM [^2] https://docs.vllm.ai/en/latest/. For each question, we generate one greedy answer with temperature = 0, and sample 20 answers with temperature = 1, top-p = 0.95, and top-k = 50. Checkpoints for all training methods are selected using the in-domain test set. All other parameters are kept at their default settings. All the prompts can be seen in §G.
@@ Appendix E Details of Baselines
In this section, we describe how each training-free baseline method is implemented. For the question $q$, suppose the greedy answer generated by the model is $\tilde{r}$ and the set of sampled answers is $\hat{\mathcal{R}}$. $\hat{\mathcal{R}}$ contains 20 responses in our paper. Using the token generation probabilities of the model to represent confidence is a common approach (Guo et al., 2017; Desai and Durrett, 2020; Jiang et al., 2021; Ni et al., 2024b); in this work, we implement two versions.
Prob.
@@
Temperature-scaling-based Methods.
Thermometer performs temperature scaling on the generation probability of the answer and adaptively assigns a temperature for each task. This is done by learning a mapping from the model’s internal states to the temperature using the training set. DACA uses a pre-trained model to calibrate the logits of a model after SFT and RLHF via a learned temperature. For both methods, we followed the approach they outlined in their paper for handling free-form QA. We append the generated answer (the same as greedy-search answer in our paper) to the prompt, asking the model to judge whether the answer is correct by choosing between two options: A. Yes and B. No. For training, we sample 1,000 examples from each dataset in HonestyBench-Train (5,000 in total). All other settings follow the original configuration.